OpenStudy (anonymous):

Last year, there were 120 seniors in Limon High School. This year there are 145. What is the percent of increase? Round to the nearest percent. A. 83% B. 17% C. 21% D. 79%

OpenStudy (anonymous):

\[\frac{ 145 - 120 }{ 120 } \times 100 \] is the increase in percentage of seniors.

OpenStudy (anonymous):

C

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

120*21%=25.2 25.2+120=145